OpenStudy (anonymous):

A car moving at 20ms-1 starts to slow down at 4ms-2 when it comes to a junction. The distance the car will travel before it comes to a stop will be ?

OpenStudy (anonymous):

use 1st equation of motion Vf=Vi+at put Vf=0 b/c the car finally come to rest. u'll get time t=5 sec now use 2nd equation of motion i.e; S=Vi(t)+1/2a(t)^2 put the values: Vi=20m/s t=5sec u'll get S=150m